# Rethinking CLIP-based Video Learners in Cross-Domain Open-Vocabulary Action Recognition

### v1

@@ 4 Methodology
In this work, we focus on tackling the domain gaps for cross-domain open-vocabulary action recognition. Specifically, we propose to address a critical challenge of the task, namely scene bias. To this end, we propose a novel Scene-Aware video-Text Alignment (SATA) method. In what follows, we illustrate our method in detail.
Figure 2: An overview of our proposed Scene-Aware video-Text Alignment (SATA) method. Our method basically includes a video encoder and a text encoder for video and text representation extraction, with a contrastive loss for video-text alignment. Based on the scene-encoded text prompts, we propose the Scene-Aware Discrimination and Scene-Aware Consistency losses, aiming to learn scene-agnostic video representations for cross-domain open-vocabulary action recognition. Best viewed in color.
@@ 4.1 Problem Formulation
In cross-domain open-vocabulary action recognition, a set of labeled videos $\mathcal{D}=\{(x,y)\}$ from a source domain are given for training, where $x$ and $y$ denote a source video and its ground-truth label index. The source domain consists of $K$ action categories, and the action names (label texts) are denoted by $\{a_{1},a_{2},\dots,a_{K}\}$. Given only source videos for training, our goal is to develop a model that is generalizable in unseen target domains, where the source and target domains follow different distributions and different label spaces. Following previous open-vocabulary action recognition works, we sample $T$ frames from each video as model input during training and test, i.e., $x=\{x_{1},x_{2},\dots,x_{T}\}$.
@@ 4.2 Model Overview
Building upon CLIP, our model consists of a video encoder $f_{\text{vid}}(\cdot)$ and a text encoder $f_{\text{txt}}(\cdot)$. Following OpenVCLIP [73], we construct a video encoder by integrating patches from neighboring frames in each self-attention layer of the original CLIP image encoder. For video $x$, the video representation $e_{x}$ is obtained by the average of local video representations, i.e., $e_{x}=f_{\text{vid}}(x)=1/T\sum_{t=1}^{T}f^{t}_{\text{vid}}(x)$, where each local representation corresponds to one frame. We use the original CLIP text encoder and keep it frozen during training. Following previous works [73, 53], we use a video-text alignment loss to adapt video data, which is formulated as follows:
@@
where $s(\cdot,\cdot)$ denotes the cosine similarity, $\tau$ is the temperature, $e_{a_{k}}=f_{\text{txt}}(g(a_{k}))$ is the text representation of the $k$-th action category. The function $g(\cdot)$ transforms an action name into a text prompt in the form of “a video of a person [doing something].”, e.g., “a video of a person abseiling.” for the action “abseiling”. To mitigate scene bias, we contribute a Scene-Aware Discrimination loss and a Scene-Aware Consistency loss for learning scene-agnostic video representations, which will be illustrated in the next subsection. An overview of our proposed SATA is shown in Figure 2.
@@ 4.3 Scene-Aware Video-Text Alignment
Table 2: Evaluation of CLIP-based video learners trained on Kinetics150 on our XOV-Action benchmark. “C”, “O” and “A” denote the accuracy (%) of closed-set, open-set and all action categories, respectively. “AVG” denotes the average performance over the four test domains. * indicates the model variant with frozen text encoder. The bold and underlined numbers indicate the best and second best.
@@ C O A C O A C O A C O A C O A
SATA (Ours) 85.53 62.01 73.71 58.70 44.78 53.79 47.35 33.54 41.81 27.55 10.97 19.04 54.78 37.82 47.09
Then, based on the scene-encoded text prompts, we design a Scene-Aware Discrimination loss, which is given as follows:
$L_{\text{dis}}=-\log{\frac{\exp{(s(e_{x},e_{a_{y}}))}}{\exp{(s(e_{x},e_{a_{y}}))}+\sum_{n=1}^{N}\exp{(s(e_{x},\tilde{e}^{n}_{a_{y}}))}}}.$ (2)
In this loss, $\tilde{e}^{n}_{a_{y}}=f_{\text{txt}}(\tilde{g}(a_{y},n))$ is the representation of a scene-encoded text prompt, which encodes semantic information of both the ground-truth action $a_{y}$ and a random scene. The function $\tilde{g}(\cdot,n)$ transforms an action name into a scene-encoded text prompt using the $n$-th scene suffix. Accordingly, our Scene-Aware Discrimination loss pushes the video representations away from the scene-encoded text representations in video-text alignment. In this way, the loss leverages the strong power of CLIP text encoder, which encourages the video encoder to pay less attention on scene information and thus pay more attention to action information, thereby bridging the domain gap.
Scene-Aware Consistency: In addition, we propose a Scene-Aware Consistency loss to constrain the video representation learning. Our intuition is that a video should have a more consistent representation with its scene-encoded text prompts compared with text representations of other action categories, since the scene-encoded text prompts encodes semantic information of the ground-truth action category (in addition to scenes). For example, consider a training video that depicts a person playing basketball in a court. The video should have a similar representation to the scene-encoded text prompts like “a video of a person playing basketball in the park.”, compared with text prompts of other action categories, e.g., “a video of a person playing football.”. Specifically, our proposed Scene-Aware Consistency loss is formulated as follows:
$L_{\text{con}}=$ $-\frac{1}{N+1}\log{\frac{\exp{(s(\hat{e}_{x},e_{a_{y}}))}}{Z}}$ (3)
$-\frac{1}{N+1}\sum\nolimits_{n=1}^{N}\log{\frac{\exp{(s(\hat{e}_{x},\tilde{e}^{n}_{a_{y}})))}}{Z}},$
where $\hat{e}_{x}=f_{\text{proj}}(f_{\text{vid}}(x))$ is the video representation after projection $f_{\text{proj}}(\cdot)$, and $Z=\sum_{k=1}^{K}\exp{(s(\hat{e}_{x},e_{a_{k}}))}+\sum_{n=1}^{N}\exp{(s(\hat{e}_{x},\tilde{e}^{n}_{a_{y}}))}$. Note that, the projection function is crucial to harmonize our Scene-Aware Discrimination and Scene-Aware Consistency losses, as shown in our experimental analysis.
Furthermore, we propose a Text-Adaptive Aggregation module to enhance video representations for classification. Specifically, we aggregate local video representations by text-adaptive weights instead of average pooling. The aggregation module is formulated as $e^{\text{txt}}_{x}=\sum_{t=1}^{T}w^{t}_{x}e^{t}_{x}$, where $e^{t}_{x}=f^{t}_{\text{vid}}(x)$ is the $t$-th local representation. And, the weight $w^{t}_{x}$ is as follows:
$w^{t}_{x}=\frac{1}{K}\sum_{k=1}^{K}\frac{\exp{(s(e^{t}_{x},e_{a_{k}})/\eta)}}{\sum_{m=1}^{T}\exp{(s(e^{m}_{x},e_{a_{k}})/\eta)}},$ (4)
where $\eta$ is the temperature. In this way, we obtain integrated video representations adaptive to specific action categories. In our implementation, we use $e^{\text{txt}}_{x}$ instead of $e_{x}$ in Eq. (1) and Eq. (2) for training, and also use $e^{\text{txt}}_{x}$ for test.
The overall training loss of our SATA is given as follows:
$L=L_{\text{vta}}+\lambda_{\text{dis}}L_{\text{dis}}+\lambda_{\text{con}}L_{\text{con}},$ (5)
where $\lambda_{\text{dis}}$ and $\lambda_{\text{con}}$ are coefficients for trade-off. During test, we use text presentations of action categories in target domains for classification.
@@
Table 3: Evaluation of CLIP-based video learners trained on Kinetics400 on our XOV-Action benchmark. We use the officially released checkpoints of previous methods for evaluation, and we use one temporal clip during inference for a fair comparison. “C”, “O” and “A” denote the ACC (%) of closed-set, open-set and all action categories, respectively. “AVG” denotes the average performance over the four test domains. * indicates the model variant with frozen text encoder. The bold and underlined numbers indicate the best and second best.
@@
SATA (Ours) 88.51 65.17 76.78 61.44 44.03 55.29 54.49 28.05 43.89 28.70 11.84 20.06 58.29 37.27 49.01
Table 4: Ablation study of our proposed SATA trained on Kinetics150. “C”, “O” and “A” denotes the accuracy (%) of closed-set, open-set and all action categories, respectively. “AVG” denotes the average performance over the four test domains. The closed-set ACC primarily evaluates the model abilities of tackling domain gaps.
@@
+$L_{\text{dis}}$ 84.63 60.43 72.47 59.41 43.10 53.65 49.18 25.31 39.61 25.93 10.53 18.03 54.79 34.84 45.94
+$L_{\text{dis}}$&$L_{\text{con}}$ 84.84 62.28 73.50 58.30 44.03 53.26 49.39 29.57 41.44 26.39 11.40 18.70 54.73 36.82 46.73
Full 85.53 62.01 73.71 58.70 44.78 53.79 47.35 33.54 41.81 27.55 10.97 19.04 54.78 37.82 47.09
+$L_{\text{dis}}$&$L_{\text{con}}$ w/o Proj 84.20 61.75 72.92 58.50 42.91 53.00 48.78 28.35 40.59 20.14 14.04 17.01 52.91 36.76 45.88
5 Experiments
5.1 Experimental Setups
Implementation details: For each video, our model takes $T=16$ frames of size $224\times 224$ as inputs. Following ViFiCLIP [53], we adopt temporal jitter, multi-scale random spatial crop and color jitter for augmentation during training. During test, we use one temporal clip composed of center-cropped frames for inference. Regarding the architecture of the video encoder, we adopt ViT-B/32 and ViT-B/16 initialized by CLIP image encoder for Kinetics150 and Kinetics400, respectively. For temporal modeling, we set the temporal receptive field of self-attention layers to 7 in the video encoder. We ask GPT-4 [47] to output 300 scene suffixes that are common in daily life, and randomly sample $N=50$ scene suffixes for Eq. (2) in each batch. The projection function $f_{\text{proj}}(\cdot)$ is an MLP with residual. The parameters are optimized using AdamW optimizer with a batch size of 512, a learning rate of 8e-6, a momentum of 0.9 and a weight decay of 1e-3. By default, we set the loss coefficients as $\lambda_{\text{dis}}=0.2$ and $\lambda_{\text{con}}=0.2$, and the temperature coefficients as $\tau=0.01$ and $\eta=0.1$. Following OpenVCLIP [73], we adopt the SWA technique [24] for producing the final model.
Evaluation protocol: We adopt five state-of-the-art CLIP-based video learners for evaluation and comparison. To conduct a solid evaluation on Kinetics400, we utilize the officially released checkpoints of each method, instead of retraining models. In the case of Kinetics150, we ensure a fair comparison by adopting consistent network architectures and training recipes across all the methods based on their official codes. During inference, we use one temporal clip for a fair comparison. We use our proposed XOV-Action benchmark for evaluation, which includes test domains of various types of domain gaps and action categories of different types. We report the accuracy of closed-set, open-set and all categories in individual test domains as well as the average accuracies, for a comprehensive analysis. When conducting a comparison between models, we prioritize the average ovarall accuracy as the primary metric.
@@ 5.2 Results
Comparison with state-of-the-arts: The results are summarized in Table 2 and 3. As shown in the tables, by adapting CLIP to video data, previous CLIP-based video learners obtain substantial improvement over the original CLIP for cross-domain open-vocabulary action recognition. Although these methods show impressive performance for the domain with a small gap (UCF), they still exhibit limited performance for the domains with large gaps (ARID and NEC-Dr). These results reveal potential challenges of the task, as discussed in Section 1 and 3. By learning scene-agnostic video representations, our proposed SATA outperforms previous state-of-the-arts by 1.69% and 1.45% in terms of average overall accuracy for Kinetics150 and Kinetics400, respectively. These improvements are significant considering that the metric represents an average across four accuracy metrics.
@@
Figure 4: Quantitative analysis of the coefficient $\lambda_{\text{dis}}$ for Scene-Aware Discrimination loss by closed-set accuracy on four test domains. The horizontal axis shows the value of $\lambda_{\text{dis}}$.
Ablation Study: Table 4 summarizes the results of our ablation study. In the table, the baseline refers to the model trained with the basic video-text alignment loss $L_{\text{vta}}$. By introducing our proposed Scene-Aware Discrimination loss $L_{\text{dis}}$, our model obtains improvement on XOV-Action in terms of closed-set ACC on all the four test domains. Specifically, our model obtains significant improvements of 1.83% and 4.63% in terms of closed-set ACC on ARID and NEC-Dr, which have large domain gaps with the training domain. This is because that our proposed Scene-Aware Discrimination loss encourages the video encoder to downweight the attention on scene information in videos and thus pay more attention to action information that are generalizable across different video domains. However, we find that introducing only the Scene-Aware Discrimination loss reduces the open-set performance. By using both Scene-Aware Discrimination loss $L_{\text{dis}}$ and Scene-Aware Consistency loss $L_{\text{con}}$, our model can obtain improved closed-set performance while maintaining the open-set performance, and this leads to an improvement of 1.07% in terms of average overall accuracy over the baseline. Notably, the projection within our Scene-Aware Consistency loss is crucial to harmonize the two losses. If the projection is dropped, the improvement is very limited as shown in the last line of Table 4. Finally, by introducing the Text-Adaptive Aggregation module, our model obtains further improvement in terms of average overall accuracy (please refer to the Appendix for more analysis of this module).
Effect of scene suffixes: In Figure 4, we show how the suffix number affects the model performance. We set the loss coefficient $\lambda_{\text{dis}}=0.5$ in this experiment. As shown in the figure, our model obtains better closed-set accuracy as the total number of scene suffixes increases. According to the design of our Scene-Aware Discrimination loss, we distinguish videos apart from more scene-encoded text prompts in feature space with more scene suffixes used. In this way, the video representations are less likely to be confused with scenes, leading to stronger abilities for recognizing actions in unseen domains.
Figure 5: Analysis of the coefficient $\lambda_{\text{con}}$ for Scene-Aware Consistency loss. The horizontal axis shows the value of $\lambda_{\text{con}}$.
@@
Effect of the Scene-Aware Discrimination loss: We conduct a quantitative analysis of the loss coefficient $\lambda_{\text{dis}}$, and the results on the four test domains are shown in Figure 4. The results in the figure show that our model consistently obtains better closed-set performance on all test domains with a larger loss weight. The comprehensive results demonstrate that our proposed Scene-Aware Discrimination loss effectively bridges the domain gaps of various types by mitigating scene bias.
Effect of the Scene-Aware Consistency loss: We conduct a quantitative analysis of the loss coefficient $\lambda_{\text{con}}$, and the results are shown in Figure 5. As shown in the figure, our model obtains higher open-set performance but lower closed-set performance when using a larger loss weight. The results empirically reveal that a good trade-off between our Scene-Aware Discrimination and Consistency losses is crucial for cross-domain open-vocabulary action recognition.
Figure 6: Qualitative analysis by attention visualization. Best viewed in color.
Qualitative analysis: In Figure 6, we show a qualitative analysis by attention visualization. As shown in the figure, compared with the baseline that pays much attention to scenes, our model focuses on the action performers when recognizing actions across domains. The results demonstrate that our model effectively mitigates scene bias.
5.3 Conclusion

### v2

@@ 4.1 Problem Formulation
In cross-domain open-vocabulary action recognition, a set of labeled videos $\mathcal{D}=\{(x,y)\}$ from a source domain are given for training, where $x$ and $y$ denote a source video and its ground-truth action label index. The source domain consists of $K$ action categories, and the action names (label texts) are denoted by $\{a_{1},a_{2},\dots,a_{K}\}$. Given only source videos for training, our goal is to develop a model that is generalizable in unseen target domains, where the source and target domains follow different distributions and different label spaces. Following previous open-vocabulary action recognition works, we sample $T$ frames from each video as model input during training and test.
Figure 2: An overview of our proposed Scene-Aware video-Text Alignment (SATA) method. Our method basically includes a video encoder and a text encoder for representation extraction, with a contrastive loss for video-text alignment. Based on the scene-encoded text prompts, we propose the Scene-Aware Discrimination and Action-Aware Discrimination losses, aiming to learn scene-agnostic video representations for cross-domain open-vocabulary action recognition. Best viewed in color.
@@ 4.2 Model Overview
Building upon CLIP, our model consists of a video encoder $f_{\text{vid}}(\cdot)$ and a text encoder $f_{\text{txt}}(\cdot)$. Following OpenVCLIP [9], we construct a video encoder by integrating patches from neighboring frames in each self-attention layer of the original CLIP image encoder. For the video $x$, the video representation $e_{x}=f_{\text{vid}}(x)$ is obtained by the average of local video representations (each local representation corresponds to one frame). We use the original CLIP text encoder and keep it frozen during training. Following previous works [9; 8], we use a video-text alignment loss to adapt video data, which is formulated as follows:
@@
where $s(\cdot,\cdot)$ denotes the cosine similarity, $\tau$ is the temperature, $e_{a_{k}}=f_{\text{txt}}(g(a_{k}))$ is the text representation of the $k$-th action category. The function $g(\cdot)$ transforms an action name into a text prompt in the form of “a video of a person [doing something].”, e.g., “a video of a person abseiling.” for the action “abseiling”. To mitigate scene bias, we contribute a Scene-Aware Discrimination loss and a Action-Aware Discrimination loss for learning scene-agnostic video representations, which will be illustrated in the next subsection. An overview of our proposed method is shown in Figure 2.
@@ 4.3 Scene-Aware Video-Text Alignment
Then, based on the scene-encoded text prompts, we design a Scene-Aware Discrimination loss, which is formulated as follows:
$L_{\text{scene}}=-\log{\frac{\exp{(s(e_{x},e_{a_{y}}))}}{\exp{(s(e_{x},e_{a_{y}}))}+\sum_{n=1}^{N}\exp{(s(e_{x},\tilde{e}^{n}_{a_{y}}))}}}.$ (2)
In this loss, $\tilde{e}^{n}_{a_{y}}=f_{\text{txt}}(\tilde{g}(a_{y},n))$ is the representation of a scene-encoded text prompt, which encodes the semantic information of the $n$-th scene. The function $\tilde{g}(\cdot,n)$ transforms an action name into a scene-encoded text prompt using the $n$-th scene suffix. According to Eq. (2), our Scene-Aware Discrimination loss pushes the video representations away from the scene-encoded text representations in video-text alignment. In this way, the loss leverages the strong power of CLIP text encoder, which encourages the video encoder to pay less attention to scene information and thus pay more attention to action information, thereby mitigating the scene bias when fitting training videos.
Action-Aware Discrimination: In our experiments, we find that although the Scene-Aware Discrimination loss $L_{\text{scene}}$ effectively improves the cross-domain action recognition performance for the closed-set categories, it reduces the cross-domain performance for open-set categories. Thus, it leads to limited improvement in the overall cross-domain open-vocabulary performance.
A critical reason is that $L_{\text{scene}}$ pushes a video away from a scene-encoded text prompt in representation space, and this may result in higher similarity between the video and some non-ground-truth action texts in representation space, relative to the scene-encoded text prompt. However, since the scene-encoded text prompt encodes semantic information of the ground-truth category (in addition to a scene), the video should have a more different representation from the non-ground-truth action texts, compared with the scene-encoded text prompt. For example, consider a video that depicts a person playing basketball in a court. This video should have a dissimilar representation to the text prompts of other categories (e.g., “a video of a person kicking soccer.”) compared with the scene-encoded text prompts (e.g., “a video of a person playing basketball in the park.”). Overall, this issue would cause some confusion in video representation space.
Accordingly, to alleviate the degradation of cross-domain open-set performance, we propose a Action-Aware Discrimination loss to constrain video representation learning. Relative to the scene-encoded text representations that encodes ground-truth action semantics, our Action-Aware Discrimination loss pushes the video representations away from the text presentations of non-ground-truth categories, which is formulated as follows:
$L_{\text{action}}=\frac{1}{N(K-1)}\sum\nolimits_{n=1}^{N}\sum\nolimits_{k\not=y}^{K}\max{\left(0,\delta-s(e_{x},\tilde{e}^{n}_{a_{y}})+s(e_{x},e_{a_{k}})\right)},\vspace{-0.1cm}$ (3)
where $\delta$ is the margin. In principle, $L_{\text{action}}$ introduces a constraint for non-ground-truth action text $a_{k}$ ($k\not=y$), i.e., $s(e_{x},\tilde{e}^{n}_{a_{y}})-s(e_{x},e_{a_{k}})\geq\delta$. By using both $L_{\text{scene}}$ and $L_{\text{action}}$, our SATA aims to learn a more reasonable video representation space, i.e., $s(e_{x},e_{a_{y}})>s(e_{x},\tilde{e}^{n}_{a_{y}})>s(e_{x},e_{a_{k}})$ when $k\not=y$.
In summary, the overall training loss of our SATA is given as follows:
$L=L_{\text{vta}}+\lambda_{\text{scene}}L_{\text{scene}}+\lambda_{\text{action}}L_{\text{action}},$ (4)
where $\lambda_{\text{scene}}$ and $\lambda_{\text{action}}$ are coefficients for trade-off. During test, we use text presentations of action categories in target domains for classification.
Table 2: Evaluation of CLIP-based video learners trained on Kinetics150 on XOV-Action. “C”, “O” and “A” denote the ACC (%) for closed-set, open-set and all action categories, respectively. “AVG” denotes the average ACC over four test domains. * indicates the model variant with frozen text encoder. The bold/underlined numbers indicate the best/second best. We prioritize the average overall ACC as the primary metric in model comparison.
@@ C O A C O A C O A C O A C O A
SATA (Ours) 84.89 62.86 73.82 58.40 44.59 53.53 48.78 29.57 41.08 24.54 12.72 18.48 54.15 37.44 46.73
5 Experiments
5.1 Experimental Setups
Implementation Details: For each video, our model takes $T=16$ frames of size $224\times 224$ as inputs. Following ViFiCLIP [8], we adopt temporal jitter, multi-scale random spatial crop and color jitter for augmentation during training. During test, we use one temporal clip composed of center-cropped frames for inference. Regarding the architecture of the video encoder, we adopt ViT-B/32 and ViT-B/16 initialized by CLIP image encoder for Kinetics150 and Kinetics400, respectively. For temporal modeling, we set the temporal receptive field of self-attention layers to 7 in the video encoder. We ask GPT-4 [99] to output 300 scene suffixes that are common in daily life, and randomly sample $N=50$ scene suffixes for our losses in each batch. The parameters are optimized using AdamW optimizer with a batch size of 512, a learning rate of 8e-6, a momentum of 0.9 and a weight decay of 1e-3. If not specified, we set the loss coefficients as $\lambda_{\text{scene}}=0.2$ and $\lambda_{\text{action}}=0.2$, the margin as $\delta=0.5$, and the temperature coefficient as $\tau=0.01$. Following OpenVCLIP [9], we adopt the SWA technique [100] for producing the final model. All experiments are conducted by PyTorch with 8 NVIDIA RTX 3090 GPUs.
@@ 5.1 Experimental Setups
Table 3: Evaluation of CLIP-based video learners trained on Kinetics400 on XOV-Action. We use the officially released checkpoints of previous methods for evaluation, and we use one temporal clip during inference for a fair comparison. “C”, “O” and “A” denote the ACC (%) for closed-set, open-set and all action categories, respectively. “AVG” denotes the average ACC over four test domains. * indicates the model variant with frozen text encoder. The bold/underlined numbers indicate the best/second best. We prioritize the average overall ACC as the primary metric in model comparison.
@@ C O A C O A C O A C O A C O A
SATA (Ours) 88.51 65.70 77.05 61.54 41.98 54.63 56.94 25.61 44.38 32.41 10.53 21.18 59.45 35.95 49.31
Table 4: Ablation study of our proposed SATA trained on Kinetics150. “C”, “O” and “A” denotes the ACC (%) for closed-set, open-set and all action categories, respectively. “AVG” denotes the average ACC over four test domains. The closed-set accuracy primarily evaluates the model abilities of tackling domain gaps when fitting training videos.
@@
+$L_{\text{scene}}$ 84.63 60.43 72.47 59.41 43.10 53.65 49.18 25.31 39.61 25.93 10.53 18.03 54.79 34.84 45.94
+$L_{\text{scene}}$&$L_{\text{action}}$ 84.89 62.86 73.82 58.40 44.59 53.53 48.78 29.57 41.08 24.54 12.72 18.48 54.15 37.44 46.73
Evaluation Protocol: We adopt five state-of-the-art CLIP-based video learners for evaluation and comparison. To conduct a solid evaluation on Kinetics400, we utilize the officially released checkpoints of each method, instead of retraining models. In the case of Kinetics150, we ensure a fair comparison by adopting consistent network architectures and training recipes across all the methods based on their official codes. During inference, we use one temporal clip for a fair comparison. We evaluate each checkpoint on all the four test domains of our XOV-Action benchmark, where different test domains have various types of domain gaps and differ in action category. For a comprehensive analysis, we report the accuracy of closed-set, open-set and all categories in individual test domains as well as the average accuracies. When conducting a comparison between different methods, we prioritize the average overall accuracy as the primary metric.
@@ 5.2 Results
Comparison with State-of-the-arts: The results are summarized in Table 2 and 3. As shown in the tables, by adapting CLIP to video data, previous CLIP-based video learners obtain substantial improvement over the original CLIP for cross-domain open-vocabulary action recognition. Although these methods show impressive performance for the domain with a small gap (UCF), they still exhibit limited performance for the domains with large gaps (ARID and NEC-Dr). These results reveal potential challenges of the task, as discussed in Section 1 and 3. By learning scene-agnostic video representations, our proposed SATA outperforms previous state-of-the-arts by 1.33% and 1.75% in terms of average overall accuracy for Kinetics150 and Kinetics400, respectively. These improvements are significant, especially considering that the metric is an average across four test domains.
Ablation Study: Table 4 summarizes the results of our ablation study. In the table, the baseline refers to the model trained with the basic video-text alignment loss $L_{\text{vta}}$. By introducing our proposed Scene-Aware Discrimination loss $L_{\text{scene}}$, our model obtains improvement on XOV-Action in terms of closed-set accuracy on all the four test domains. Specifically, our model obtains significant improvements of 1.83% and 4.63% in terms of closed-set accuracy on ARID and NEC-Dr, which have large domain gaps with the training domain. This is because that $L_{\text{scene}}$ encourages the video encoder to downweight the attention on scene information in videos and thus pay more attention to action information that are generalizable across different domains. However, we find that introducing only the Scene-Aware Discrimination loss reduces the open-set performance, leading to limited improvement in average overall accuracy on XOV-Action. By further introducing the Action-Aware Discrimination loss $L_{\text{action}}$, our model can obtain improved cross-domain closed-set accuracy over the baseline while maintaining the open-set accuracy on each test domain as much as possible, and this leads to an improvement of 1.07% in terms of average overall accuracy over the baseline.
@@
Figure 4: Quantitative analysis of the coefficient $\lambda_{\text{scene}}$ for the Scene-Aware Discrimination loss by closed-set accuracy on four test domains. The horizontal axis shows the value of $\lambda_{\text{scene}}$.
@@
Effect of Scene-Aware Discrimination: We conduct a quantitative analysis to the loss coefficient $\lambda_{\text{scene}}$, and the results on the four test domains are shown in Figure 4. The results in the figure show that our model consistently obtains better closed-set performance on all test domains with a larger loss weight. The comprehensive results demonstrate that our proposed Scene-Aware Discrimination loss effectively tackles various types of domain gaps by mitigating scene bias in fitting training videos.
Effect of Scene Suffixes: In Figure 4, we show how the suffix number affects the model performance. To show the loss effect more clearly, we set the loss coefficient $\lambda_{\text{scene}}=0.5$ in this experiment. As shown in the figure, our model obtains better closed-set accuracy as the total number of scene suffixes increases. According to the design of our Scene-Aware Discrimination loss, we distinguish
Figure 5: Qualitative analysis by t-SNE [101].
videos apart from more scene-encoded text prompts in representation space with more scene suffixes used. In this way, the video representations are less likely to be confused with scenes, leading to stronger abilities for recognizing actions in unseen domains.
Effect of Action-Aware Discrimination: Figure 5 intuitively shows how the Action-Aware Discrimination loss $L_{\text{action}}$ affects representation space. As shown in the figure, without the $L_{\text{action}}$ introduced, our Scene-Aware Discrimination loss may lead to higher similarity between a video and some non-ground-truth action texts compared with the scene-encoded text prompts in representation space, which would lead to confusion in video classification. By introducing $L_{\text{action}}$, our SATA produces a much more reasonable representation space, where a video is more distant from the non-ground-truth action texts compared with the scene-encoded text prompts. For the analysis to the loss weight $\lambda_{\text{action}}$, please refer to the Appendix.
6 Conclusion